OpenStudy (anonymous):

A force of 20 pounds stretches a spring 3/4 foot on anexercise machine. Find the work done in stretching the spring 1 foot.

OpenStudy (zarkon):

use Hooke's law integrate

OpenStudy (anonymous):

ShoW me how, please. There's a reason why I'm asking

OpenStudy (zarkon):

you know hooke's law?

OpenStudy (anonymous):

Not well

OpenStudy (zarkon):

\[F=-kx\] F=20 x=3/4

OpenStudy (zarkon):

what is k?

OpenStudy (zarkon):

then if you are stretching from its natural length to 1 foot beyond then the work done is \[\int_{0}^{1}(-kx)dx\]

OpenStudy (anonymous):

So would it be 15ft-lb?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

20=-k3/4 \[k=-20\frac{4}{3}\] \[k=\frac{-80}{3}\]

OpenStudy (zarkon):

now integrate

OpenStudy (anonymous):

Shouldn't I leave it as 80/3ft-lb

OpenStudy (zarkon):

keep it as \[-\frac{80}{3}\] then integrate

OpenStudy (zarkon):

you need to calculate the work.. \[\int_{0}^{1}(-kx)dx=\int_{0}^{1}\left(-\frac{-80}{3}x\right)dx=\int_{0}^{1}\left(\frac{80}{3}x\right)dx\]

OpenStudy (anonymous):

Final answer is supposed to be in that form. My only other option is either 40/3 or 80/9. Can't seem to get either

OpenStudy (zarkon):

it is 40/3ft-lb \[\int_{0}^{1}\left(\frac{80}{3}x\right)dx=\left.\frac{80}{3}\frac{x^2}{2}\right|_{0}^{1}\] \[=\frac{40}{3}1^2-\frac{40}{3}0^2=\frac{40}{3}ft-lb\]